Help me to solve this question
mike and ryan have $ 260 altogether. after mike spend 1/4 of his money and ryan spend $ 64, now they have same amount of money. How much money each of them have at first ?

Answer by stanbon(75887)

mike and ryan have $ 260 altogether. after mike spend 1/4 of his money and ryan spend $ 64, now they have same amount of money. How much money each of them have at first ?
----
Equations:
m + r = 260 dollars
(m-(1/4)m) = (r-64)
-------------------
Rearrange:
r = 260-m
(3/4)m = r-64
----
Substitute for "r" and solve for "m":
(3/4)m = 260-m-64
(7/4)m = 196
m = (4/7)196
m = $112 (mike's original amount)
---
solve for "r":
r = 260 -m
r = 260 - 112
r = $148 (ryan's orinal amount)
===================================
Cheers,
Stan H.
============